OpenStudy (anonymous):

Write an equation for the line that passes through the following point: (-4, -2), (2,6) Explain

Directrix (directrix):

Do you know how to get the slope determined by these two points?

OpenStudy (anonymous):

y2-y1/x2-x1 I'm guessing?

OpenStudy (anonymous):

So y2-y1/x2-x1 is correct?

Directrix (directrix):

slope is rise over run which is what you have (y2-y1)/(x2-x1) . Think of it as change in y over change in x. So, now, crank out the value of m, the slope. Either point can be designated as (x1, y1).

Directrix (directrix):

Yes, with the ( ). Otherwise, it could be read as 10 - 6/6 - 4 which is 10 - 1 - 4 = 5. When, you really wanted (10-6)/(6-4) = 4/2 = 2. I just made up those numbers to illustrate that grouping symbols matter.

Directrix (directrix):

If you were writing horizontally, it wouldn't be problem. ------------------- Compute the slope now and post here.

OpenStudy (anonymous):

For slope I got 4/3. But for the rest, I don't know what to do.

Directrix (directrix):

Be glad you got that right and don't worry about what you can't do next.

Directrix (directrix):

We'll use the point/slope form of a line. Actually, all these forms of writing a line amount to the same thing in the end.

Directrix (directrix):

y - y1 = m (x - x1) where m is the slope and (x1,y1) is either of the two points. Two points determine a unique line so it doesn't matter which one you use.

Directrix (directrix):

I'll us (-4,-2) Y - (-2) = (4/3) [ x - (-4)] It is extremely important to use grouping symbols here because it is so easy to get the minus symbol for subtraction mixed up with the negative sign of the number. y + 2 = (4/3) [ x + 4] See if you agree so far.

OpenStudy (amistre64):

2 negatives makes a positive :)

Directrix (directrix):

@sakigirl --> Somebody is here in your place answering your questions.

OpenStudy (anonymous):

Oh! It's coming back to me now :) Thanks!

Directrix (directrix):

Where did you go? I thought we were having a dialogue. I'm waiting for you to answer this question. y + 2 = (4/3) [ x + 4] See if you agree so far.

OpenStudy (anonymous):

agree

Directrix (directrix):

It's a good idea to stick with one problem rather than to try to do several problems at once. A person (amistre64) above you wrote the following: 2 negatives makes a positive :) How is that statement relevant to the problem? Is there an error so far?

OpenStudy (anonymous):

-+-=+

OpenStudy (anonymous):

No error so far

Directrix (directrix):

y + 2 = (4/3) [ x + 4] y + 2 = (4/3)x + 16/3)

Directrix (directrix):

Multiply through by 3.

Directrix (directrix):

Write what results from that here.

OpenStudy (anonymous):

I'm a bit confused on how you got 16/3

Directrix (directrix):

4/3 * 4 = 4/3 * 4/1 = 16/3

Directrix (directrix):

Your task is to multiply y + 2 = (4/3)x + 16/3) by 3. Do you want to do this problem in a different way?

OpenStudy (anonymous):

It's just y=4/3x

OpenStudy (anonymous):

?

Directrix (directrix):

What about it? It's not 4 over 3x. It is (4/3) times x. 3 (4/3) x = 3/1 * 4/3 * x/1 = 4x.

OpenStudy (anonymous):

I didn't multiply it by 3 and I got y=(4/3)x.

Directrix (directrix):

y + 2 = 4/3 (x + 4) 3y + 6 = 4x + 16 3y = 4x + 16 - 6 3y = 4x + 10 y = (4/3)x + 10/3

OpenStudy (anonymous):

Oh never mind. I see where I made my mistake. Thanks!

Directrix (directrix):

This is not the only way to do the problem.

OpenStudy (anonymous):

Oh, but this is the only way I learned it since 8th grade so I prefer this way :P